Which is greater in each of the following:
-5/6, -4/3

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Dear Student

The LCM of the denominators 6 and 3 is 6
∴(-5/6) = [(-5 × 1)/ (6 × 1)] = (-5/6)
And (-4/3) = [(-4 × 2)/ (3 × 2)] = (-12/6)
Now,
-5 > -12
So, (-5/6) > (- 12/6)
∴-5/6 > -12/6
Hence, - 5/6 is greater.

Thanks